Grade 12Algebra

Find the product, using suitable properties:
26 × (– 48) + (– 48) × (–36)

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

The given equation is in the form of Distributive law of Multiplication over Addition.
= a × (b + c) = (a × b) + (a × c)
Let, a = -48, b = 26, c = -36
Now,
= 26 × (–48) + (–48) × (–36)
= -48 × (26 + (-36)
= -48 × (26–36)
= -48 × (-10)
= 480 ... [∵(- × - = +)

Thanks